# Time-Series Foundation Model for Value-at-Risk Forecasting

### v6

@@ Abstract
This study is the first to analyze the performance of a time-series foundation model for Value-at-Risk (VaR), which essentially forecasts the left-tail quantiles of returns. Foundation models, pre-trained on diverse datasets, can be applied in a zero-shot setting with minimal data or further improved through finetuning. We compare Google’s TimesFM model to conventional parametric and non-parametric models, including GARCH and Generalized Autoregressive Score (GAS), using 19 years of daily returns from the S&P 100 index and its constituents. Backtesting with over 8.5 years of out-of-sample data shows that the fine-tuned foundation model consistently outperforms traditional methods in actual-over-expected ratios. For the quantile score loss function, it performs comparably to the best econometric model, GAS. Overall, the foundation model ranks as the best or among the top performers across the 0.01, 0.025, 0.05, and 0.1 quantile forecasting. Fine-tuning significantly improves accuracy, showing that zero-shot use is not optimal for VaR.
@@
Time-Series Forecasting , AI , Value-at-Risk Forecasting , Quantiles , TimesFM , Time series , Foundation Models , Econometric models
journal: International Journal of Forecasting
[…]
[inst1]organization=Tampere University, Department of Computing Sciences, Financial Computing and Data Analytics Group,addressline=Kalevantie 4, city=Tampere, postcode=33100, country=Finland
[…]
[inst2]organization=IIT Ropar, Department of Mathematics,addressline=Ropar, city=Punjab, postcode=140001, country=India
@@ 1 Introduction
While Machine Learning (ML) has significantly advanced financial market research—such as in asset pricing, volatility forecasting, limit order book dynamics, and optimal hedging (Gu et al., 2020; Christensen et al., 2023; Tran et al., 2018; Mikkilä and Kanniainen, 2023)—the risk management literature still largely relies on traditional econometric techniques in time-series forcasting, especially in the area of Value-at-Risk (VaR). VaR is essentially about forecasting the left-tail quantiles (e.g., the 5th or 1st percentile) of return distributions to quantify potential losses over a given time period. VaR forecasts are often based on daily return observations (for example, see González-Rivera et al., 2004; Patton et al., 2019; Taylor, 2019; Le, 2020; Taylor, 2020, based solely on historical daily returns), and even 20 years of data would yield only around 5,000 observations. This is quite limited, especially considering that DL models often contain millions of parameters that need to be learned.
@@
However, a solution to this problem is currently underway. Recently, ML has experienced a paradigm shift with the emergence of foundation models, which are large-scale, general-purpose pre-trained and often fine-tunable models on diverse datasets spanning various data distributions. Although foundation models are based on deep learning and transfer learning principles, their scale leads to the emergence of new capabilities (Bommasani et al., 2021). In 2023–2024, following major advancements in large language models (LLMs), several pre-trained time-series foundation models have emerged, including Google’s TimesFM (Das et al., 2024), Time-GPT (Garza and Mergenthaler-Canseco, 2024), and LagLlama (Rasul et al., 2023).
@@
AI-driven advancements in financial risk modeling have opened doors to new methodologies for assessing complex financial data (Tran et al., 2019; Guijarro-Ordonez et al., 2021; Chen et al., 2024). Historically, methods like GARCH-based models have been effective for capturing dependencies and volatility clustering in market data (Hoga and Demetrescu, 2023). Recent approaches, however, such as Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), and Conditional GANs (CGANs), enhance distribution forecasting by generating realistic synthetic data (Nareklishvili et al., 2023; Athey et al., 2021; Hofert et al., 2022). Such models provide flexible, non-parametric approaches for data structures that traditional econometric tools may struggle to approximate. Ericson et al. (2024) presents a comparative review showing that deep generative methods outperform traditional benchmarks in modeling complex financial dynamics for VaR forecasting. These advancements highlight the potential of AI-driven approaches in risk forecasting. This study stands out from previous research by adopting a time-series foundation model, pre-trained on diverse datasets, to forecast VaR. Unlike traditional supervised ML models trained for specific tasks, time-series foundation models are pre-trained on vast datasets spanning billions of time points. A single pre-trained model can be applied across a wide range of domains, from traffic forecasting to stock price predictions, offering quick and cost-effective solutions. The term ‘foundational’ emphasizes their flexibility, as they can be adapted for different tasks or domains with minimal additional training. The foundation models can be used with or without fine-tuning.
@@
We implement a pre-trained time-series foundation model, specifically Google’s TimesFM, for VaR forecasting and compare it with existing state-of-the-art parametric and semi-parametric approaches, including GARCH model, Generalized Autoregressive Score (GAS) one-factor model, and empirical quantile estimates. The key question is whether the pre-trained foundation model yields results comparable to, or even capable of outperforming, existing econometric methods. An additional objective is to assess the extent to which fine-tuning improves the accuracy of VaR predictions. While several foundation models have been recently introduced, this study focuses on TimesFM due to its open-source availability [^1] https://github.com/google-research/TimesFM., modifiability, and free accessibility, ensuring the results are fully replicable. [^2] The codes of this paper can be found at https://github.com/Anubha0812/TimesFM-for-Value-at-Risk. Furthermore, the authors of TimesFM have empirically demonstrated that their model outperforms TimeGPT in forecasting accuracy, adding further motivation for its selection. Rather than identifying the best foundation model, which is subject to frequent updates and new developments, this study aims to address the fundamental question of whether state-of-the-art parametric and semi-parametric methods can be outperformed.
@@
We conduct an extensive empirical evaluation of the performance of the foundation time-series model using daily return data on the 91 constituents of the S&P 100 index over 19 years from January 2005 September 2023. The first ten years of data is used for estimating (learning) model parameters and the remaining 9 years for out-of-sample testing the performance of various models. For each constituent, we calculate 1-day VaR on a daily basis. The foundation model is applied both with and without fine-tuning to understand how much the results improve with fine-tuning, which has certain computational cost. The fine-tuning is performed on a daily basis using a rolling window technique.
@@ 2.1 Time-Series Foundation Models
TimesFM model (Das et al., 2024) is a decoder-only transformer architecture, similar to GPT models. In contrast to encoder-decoder models, where the encoder processes the input sequence and the decoder generates output based on the encoded input, decoder-only models focus solely on output generation processing the input through successive decoder layers. This allows for faster training and inference times compared to encoder-decoder models. Decoder-only models like are highly advantageous for zero-shot forecasting, where minimal fine-tuning is required. On the other hand, the advantage of the encoder-decoder architecture, which is used, for example, in TimeGPT (Garza and Mergenthaler-Canseco, 2024), is that it can handle more complex inputs and exogenous variables. However, in this paper, we use univariate and well-structured time-series data on returns, and VaR is calculated separately for each security, and for that reason, the encoder component does not add significant value.
The architecture of TimesFM is optimized for long-horizon forecasting through the use of a patching strategy, where the time-series data is broken into patches. These patches function similarly to tokens in language models, significantly improving efficiency. Unlike some autoregressive models that predict one step at a time, TimesFM predicts longer output patches at once, reducing the number of necessary auto-regressive steps. This makes the model particularly well-suited for handling varying context lengths, prediction lengths, and temporal granularities during inference.
In pretraining, TimesFM was trained on a vast corpus comprising both real-world time-series data such as Google trends, Wiki Pageviews, Synthetic Data, and Electricity, Traffic, and Weather datasets used in M4 competition (Makridakis et al., 2020; Spiliotis et al., 2020). However, financial data was not used, and this aspect makes the application of the present paper particularly interesting. This corpus, although smaller in scale compared to text-based models, spans around 100 billion time points, providing the diversity needed for the model to generalize well across domains and forecasting tasks. Compared to the state-of-the-art LLMs, TimesFM time-series foundation model is much smaller, with 200M parameters and pretraining data consisting of around 100B timepoints. Despite its smaller scale, it was reported to achieve relatively high zero-shot forecasting accuracy.
@@ 2.2 Benchmark Models for VaR
  • 2.
    ARMA-GARCH Models (Angelidis et al., 2004):
@@
    3. (c)
      Empirical Distribution Function: As a nonparametric alternative, we use the Empirical Distribution Function (EDF) method (Manganelli and Engle, 2004). This approach estimates the distribution of standardized residuals $\eta_{t}$ from historical data, avoiding any parametric assumptions.
  • 3.
    One factor Gas model (Patton et al., 2019): This model belongs to the class of CAViaR models (Taylor, 2005), where the evolution of the risk measure is directly parameterized without the need to specify a full conditional distribution for asset returns. [^3] GAS model was originally proposed in Creal et al. (2013). Patton et al. (2019) proposed a dynamic GAS model employing the class of loss functions proposed in Fissler and Ziegel (2016), called FZ loss functions. In this paper, we adopt the GAS model from Section 2.3 in Patton et al. (2019). The evolution of VaR for the Gas 1 factor model is given
@@
    where $b<a<0<\gamma$ and we set $\omega=0$ as in Patton et al. (2019)
@@ 3.1 Experiment Setup
For GARCH and GAS models, we followed Patton et al. (2019) and used January 2005 to December 2014 as a training period for estimating model parameters, which were then used to forecast over the out-of-sample test period from the 2nd January 2015 to the 19th September 2023, thus resulting in a sequence of 2,268 VaR forecasts. For the rolling window technique with the Historical Method, we used the data of 512 days preceding the first test day (i.e., 2nd January 2015), and forecasted the VaR at different levels. We then shift the training sample by one day forward in time to include more recent data and exclude the oldest data point such that a fixed size of the rolling window is maintained. At each rolling step, we obtain the prediction for the next day using the most recent 512 observations, thus resulting in a sequence of 2,268 VaR forecasts for each quantile level.
@@
To better align the model with our objective, we conducted a fine-tuning procedure using the stock returns of S&P100 and its 91 constituents. We adopted the code “finetuning” in TimesFM library. More specifically, fine-tuning was carried out using a PatchedDecoderFinetuneModel with specific adjustments tailored for our task. During fine-tuning, we applied linear probing where only the core layer parameters were updated but the transformer layers were kept fixed. To prevent gradient explosion, we applied gradient clipping with a threshold of 100. The optimization process employed the Adam optimizer, that followed a cosine schedule for learning rate that started at $1\times 10^{-3}$ and decayed to $1\times 10^{-4}$ over 40,000 steps. Further, an exponential moving average decay of 0.9999 was also included in the optimization. The training loop ran for up to 100 epochs, with a provision of early stopping tied to the loss on the validation set and also a patience level was set to 5 epochs without improvement. Throughout the fine-tuning procedure, check-points were saved and finally the best model was restored for evaluation purposes.
@@ 4 Evaluation metrics
where $r_{t}$ is the realized return at time $t$. Thus, $A=\sum_{t=1}^{T}I_{t}$ denotes the number of violations in the sample. From Kupiec (1995), $A$ follows a binomial distribution, i.e., $A\sim B(T,\alpha)$, thus, under the null that $A/T=\alpha$, the relevant likelihood ratio statistic is
@@
As stated in Kupiec (1995), the UC test is capable of rejecting a model for both low and high failures, but has poor power in general. Further, the UC test tests the number of violations, while it is crucial to test whether these losses are spread evenly over time or not. The Conditional Coverage (CC) test, proposed by Christoffersen (1998), does the same, that is, it tests whether the violations are randomly distributed over time and not clustered. That is, CC test takes into account any conditionality in VaR forecasts: the model must respond to clustering events such as if volatilities are high in some periods and low in others. The test statistics is given by
@@
In the evaluation of VaR forecasts, the quantile score is a commonly used metric and is now standard in the VaR literature (Ehm et al., 2016; Gneiting, 2011; Taylor, 2019; Bams et al., 2017). The above mentioned statistical tests are good to evaluate the statistical adequacy of VaR forecasts, that is, a model for VaR forecasting is deemed to be an adequate model if we fail to reject the null hypothesis. However, these tests cannot conclude whether an ‘adequate’ model outperforms another ‘adequate’ model. Scoring functions allow us to do an honest assessment of VaR as a perfect VaR estimate would always minimize the scoring function. The quantile score $QS_{t}$ can be defined as follows:
@@
For every instant $t$, a model’s score increases either by (a) $\alpha-1$ fraction of the distance between the realized return and the VaR forecast, if realized return is smaller than the VaR forecast for the same period, or (b) by $\alpha$ fraction of the distance otherwise. This scoring rule is consistent for quantile estimation, meaning that it achieves its minimum value when the quantile forecast $q_{t}$ correctly estimates the true quantile of the distribution. Therefore, models that produce lower average quantile scores in the out-of-sample period are considered superior. Finally, to compare the performance of two models, we use the Diebold–Mariano (DM) test (Diebold and Mariano, 2002) with

### v7

# Time-Series Foundation AI Model for Value-at-Risk Forecasting

@@ Abstract
This study is the first to analyze the performance of a time-series foundation AI model for Value-at-Risk (VaR), which essentially forecasts the left-tail quantiles of returns. Foundation models, pre-trained on diverse datasets, can be applied in a zero-shot setting with minimal data or further improved through finetuning. We compare Google’s TimesFM model to conventional parametric and non-parametric models, including GARCH and Generalized Autoregressive Score (GAS), using 19 years of daily returns from the S&P 100 index and its constituents. Backtesting with over 8.5 years of out-of-sample data shows that the fine-tuned foundation model consistently outperforms traditional methods in actual-over-expected ratios. For the quantile score loss function, it performs comparably to the best econometric model, GAS. Overall, the foundation model ranks as the best or among the top performers across the 0.01, 0.025, 0.05, and 0.1 quantile forecasting. Fine-tuning significantly improves accuracy, showing that zero-shot use is not optimal for VaR.
@@
Time-Series Forecasting, AI, Value-at-Risk Forecasting, TimesFM, Time Series, Foundation Models, Econometric Models
@@ 1 Introduction
Although machine learning (ML) has significantly advanced financial market research—such as asset pricing, volatility forecasting, limit order book dynamics, and optimal hedging [1, 2, 3, 4], the risk management literature still largely relies on traditional econometric techniques in time series forecasting, especially in the area of Value-at-Risk (VaR). VaR is essentially about forecasting the left-tail quantiles (e.g., the 5th or 1st percentile) of return distributions to quantify potential losses over a given time period. VaR forecasts are often based on daily return observations [for example, see 5, 6, 7, 8, 9, 10, based solely on historical daily returns], and even 20 years of data would yield only around 5,000 observations. This is quite limited, especially considering that DL models often contain millions of parameters that need to be learned.
@@
However, a solution to this problem is currently underway. Recently, ML has experienced a paradigm shift with the emergence of foundation models, which are large-scale, general-purpose pretrained, and often fine-tunable models on diverse datasets spanning various data distributions. Although the foundation models are based on deep learning and transfer learning principles, their scale leads to the emergence of new capabilities [11]. In 2023–2024, following major advancements in large language models (LLMs), several pre-trained time-series foundation models have emerged, including Google’s TimesFM [12], Time-GPT [13], and LagLlama [14].
@@
AI-driven advancements in financial risk modeling have opened doors to new methodologies for assessing complex financial data [15, 16, 17]. Historically, methods such as GARCH-based models have been effective in capturing dependencies and clustering of volatility in market data [18]. However, recent approaches, such as Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), and Conditional GANs (CGANs), enhance distribution forecasting by generating realistic synthetic data [19, 20, 21]. Such models provide flexible, non-parametric approaches for data structures that traditional econometric tools may struggle to approximate. [22] presents a comparative review showing that deep-generative methods outperform traditional benchmarks in modeling complex financial dynamics for VaR forecasting. These advancements highlight the potential of AI-driven approaches in risk forecasting. This study stands out from previous research by adopting a time-series foundation model, pre-trained on diverse datasets, to forecast VaR. Unlike traditional supervised ML models trained for specific tasks, time-series foundation models are pre-trained on vast datasets spanning billions of time points. A single pre-trained model can be applied across a wide range of domains, from traffic forecasting to stock price predictions, offering quick and cost-effective solutions. The term ‘foundational’ emphasizes their flexibility, as they can be adapted for different tasks or domains with minimal additional training. The foundation models can be used with or without fine-tuning.
@@
We implement a pre-trained time-series foundation model, specifically Google’s TimesFM [12], for VaR forecasting and compare it with existing state-of-the-art parametric and semi-parametric approaches, including GARCH model, Generalized Autoregressive Score (GAS) one-factor model, and empirical quantile estimates [5]. The key question is whether the pre-trained foundation model yields results comparable to, or even capable of outperforming, existing econometric methods. An additional objective is to assess the extent to which fine-tuning improves the accuracy of VaR predictions.
Rather than identifying the best foundation model, which is subject to frequent updates and new developments, this study aims to address the fundamental question of whether state-of-the-art parametric and semi-parametric methods can be outperformed. While several foundation models have been recently introduced, this study focuses on TimesFM due to its robust performance, flexibility, and openness, which align well with the methodological and practical needs of our research. Other contributing factors in the selection include its open-source availability [^1] https://github.com/google-research/TimesFM., modifiability, and free accessibility, ensuring the results are fully replicable. [^2] The codes of this paper can be found at https://github.com/Anubha0812/TimesFM-for-Value-at-Risk. To be more precise, unlike proprietary models such as TimeGPT, TimesFM is open and fine-tuneable, allowing researchers to adapt the model to domain-specific datasets. TimesFM is pre-trained on a large and diverse collection of public and synthetic time-series datasets, enabling it to deliver strong zero-shot performance across a variety of forecasting tasks. One of TimesFM’s key advantages is its flexibility at inference time: it supports varying input lengths, forecast horizons, and time granularities, making it highly adaptable across different forecasting scenarios. TimesFM provides probabilistic forecasts, delivering quantile estimates (deciles) along with the mean forecast. These quantile heads can be fine-tuned as well making it more suitable for the current study. Furthermore, the authors of TimesFM have empirically demonstrated that their model outperforms TimeGPT in forecasting accuracy, adding further motivation for its selection. Finally, TimesFM has been integrated into GIFT-Eval  [^3] https://huggingface.co/spaces/Salesforce/GIFT-Eval, a comprehensive benchmarking framework for time-series models. Within this framework, it ranks at the top based on aggregated Mean Absolute Scaled Error (MASE) and Continuous Ranked Probability Score (CRPS), demonstrating its superior generalization capabilities.
@@
We conduct an extensive empirical evaluation of the performance of the foundation time-series model using daily return data on the 91 constituents of the S&P 100 index over 19 years from January 2005 September 2023. The first ten years of data is used for estimating (learning) model parameters, and the remaining 9 years for out-of-sample testing the performance of various models. For each constituent, we calculate 1-day VaR on a daily basis. The foundation model is applied both with and without fine-tuning to understand how much the results improve with fine-tuning, which has certain computational cost. The fine-tuning is performed on a daily basis using a rolling window technique.
@@ 2.1 Time-Series Foundation Models
TimesFM model [12] is a decoder-only transformer architecture, similar to GPT models. In contrast to encoder-decoder models, where the encoder processes the input sequence and the decoder generates output based on the encoded input, decoder-only models focus solely on output generation, processing the input through successive decoder layers. This allows for faster training and inference times compared to encoder-decoder models. Decoder-only models like are highly advantageous for zero-shot forecasting, where minimal fine-tuning is required. On the other hand, the advantage of the encoder-decoder architecture, which is used, for example, in TimeGPT [13], is that it can handle more complex inputs and exogenous variables. However, in this paper, we use univariate and well-structured time-series data on returns, and VaR is calculated separately for each security, and for that reason, the encoder component does not add significant value.
Figure 1: TimesFM structure Note. From [12]. A Decoder-Only Foundation Model for Time-Series Forecasting. Proceedings of the Forty-First International Conference on Machine Learning (ICML). Copyright 2024 by the Authors. The figure presents the architecture of TimesFM, a time-series foundation model built on a decoder-only transformer framework.
The architecture of TimesFM is optimized for long-horizon forecasting through a decoder-only transformer framework that employs a strategic patching mechanism. Time-series data is segmented into non-overlapping patches, which serve as analogues to tokens in language models. This patch-based representation not only improves computational efficiency by reducing input sequence length, but also enhances the model’s capacity to capture temporal dependencies over extended horizons. Each input patch is transformed via a residual MLP block and enriched with positional encodings before being processed by stacked multi-head causal self-attention layers. Unlike traditional autoregressive models that forecast one step at a time, TimesFM predicts longer output patches in parallel, significantly reducing the number of autoregressive iterations needed during inference. This architectural design makes the model highly adaptable to varying context lengths, forecasting horizons, and temporal granularities. Furthermore, the use of dynamic patch masking during training ensures that the model generalizes well across different time series patterns and lengths, making it a robust choice for zero-shot forecasting tasks in financial applications (see Figure 1 for a schematic overview of the TimesFM architecture).
In pretraining, TimesFM was trained on a vast corpus comprising both real-world time-series data such as Google trends, Wiki Pageviews, Synthetic Data, and Electricity, Traffic, and Weather datasets used in M4 competition [23, 24]. However, financial data was not used, and this aspect makes the application of the present paper particularly interesting. This corpus, although smaller in scale compared to text-based models, spans around 100 billion time points, providing the diversity needed for the model to generalize well across domains and forecasting tasks. Compared to the state-of-the-art LLMs, TimesFM time-series foundation model is much smaller, with 200M parameters and pretraining data consisting of around 100B timepoints. Despite its smaller scale, it was reported to achieve relatively high zero-shot forecasting accuracy.
@@ 2.2 Benchmark Models for VaR
  • ARMA-GARCH Models [25]:
@@
    3. 3.
      Empirical Distribution Function: As a nonparametric alternative, we use the Empirical Distribution Function (EDF) method [26]. This approach estimates the distribution of standardized residuals $\eta_{t}$ from historical data, avoiding any parametric assumptions.
  • One factor Gas model [6]: This model belongs to the class of CAViaR models [27], where the evolution of the risk measure is directly parameterized without the need to specify a full conditional distribution for asset returns. [^4] GAS model was originally proposed in [28]. [6] proposed a dynamic GAS model employing the class of loss functions proposed in [29], called FZ loss functions. In this paper, we adopt the GAS model from Section 2.3 in [6]. The evolution of VaR for the Gas 1 factor model is given
@@
    where $b<a<0<\gamma$ and we set $\omega=0$ as in [6]
@@ 3.1 Experiment Setup
For GARCH and GAS models, we followed [6] and used January 2005 to December 2014 as a training period for estimating model parameters, which were then used to forecast over the out-of-sample test period from the 2nd January 2015 to the 19th September 2023, thus resulting in a sequence of 2,268 VaR forecasts. For the rolling window technique with the Historical Method, we used the data of 512 days preceding the first test day (i.e., 2nd January 2015), and forecasted the VaR at different levels. We then shift the training sample by one day forward in time to include more recent data and exclude the oldest data point such that a fixed size of the rolling window is maintained. At each rolling step, we obtain the prediction for the next day using the most recent 512 observations, thus resulting in a sequence of 2,268 VaR forecasts for each quantile level.
@@
To better align the model with our objective, we conducted a fine-tuning procedure using the stock returns of S&P100 and its 91 constituents. We adopted the code “finetuning" in TimesFM library. More specifically, fine-tuning was carried out using a PatchedDecoderFinetuneModel with specific adjustments tailored for our task. During fine-tuning, we applied linear probing where only the core layer parameters were updated but the transformer layers were kept fixed. To prevent gradient explosion, we applied gradient clipping with a threshold of 100. The optimization process employed the Adam optimizer, that followed a cosine schedule for learning rate that started at $1\times 10^{-3}$ and decayed to $1\times 10^{-4}$ over 40,000 steps. Further, an exponential moving average decay of 0.9999 was also included in the optimization. The training loop ran for up to 100 epochs, with a provision of early stopping tied to the loss on the validation set and also a patience level was set to 5 epochs without improvement. Throughout the fine-tuning procedure, check-points were saved and finally the best model was restored for evaluation purposes.
@@ 4 Evaluation metrics
where $r_{t}$ is the realized return at time $t$. Thus, $A=\sum_{t=1}^{T}I_{t}$ denotes the number of violations in the sample. From [30], $A$ follows a binomial distribution, i.e., $A\sim B(T,\alpha)$, thus, under the null that $A/T=\alpha$, the relevant likelihood ratio statistic is
@@
As stated in [30], the UC test is capable of rejecting a model for both low and high failures, but has poor power in general. Further, the UC test tests the number of violations, while it is crucial to test whether these losses are spread evenly over time or not. The Conditional Coverage (CC) test, proposed by [31], does the same, that is, it tests whether the violations are randomly distributed over time and not clustered. That is, CC test takes into account any conditionality in VaR forecasts: the model must respond to clustering events such as if volatilities are high in some periods and low in others. The test statistics is given by
@@
In the evaluation of VaR forecasts, the quantile score is a commonly used metric and is now standard in the VaR literature [32, 33, 7, 34]. The above mentioned statistical tests are good to evaluate the statistical adequacy of VaR forecasts, that is, a model for VaR forecasting is deemed to be an adequate model if we fail to reject the null hypothesis. However, these tests cannot conclude whether an ‘adequate’ model outperforms another ‘adequate’ model. Scoring functions allow us to do an honest assessment of VaR as a perfect VaR estimate would always minimize the scoring function. The quantile score $QS_{t}$ can be defined as follows:
@@
For every instant $t$, a model’s score increases either by (a) $\alpha-1$ fraction of the distance between the realized return and the VaR forecast, if realized return is smaller than the VaR forecast for the same period, or (b) by $\alpha$ fraction of the distance otherwise. This scoring rule is consistent for quantile estimation, meaning that it achieves its minimum value when the quantile forecast $q_{t}$ correctly estimates the true quantile of the distribution. Therefore, models that produce lower average quantile scores in the out-of-sample period are considered superior. Finally, to compare the performance of two models, we use the Diebold–Mariano (DM) test [35] with